# Modified Debye-Hückel-Onsager Theory for Electrical Conductivity in Aqueous Electrolyte Solutions. Nonlocal Electrostatics Approach

### v1

@@ IV Electrophoretic contribution
$m_{a}^{(e)}=-\frac{2\kappa^{2}}{3\eta}\int\frac{d\mathbb{k}}{(2\pi)^{3}}\frac{|\varrho_{\mathbb{k}}|^{2}}{\mathbb{k}^{2}(\kappa^{2}|\varrho_{\mathbb{k}}|^{2}+\mathbb{k}^{2})},$ (56)
so that the electrophoretic contribution to the conductivity
$\Delta\sigma_{e}=\sum\limits_{a}q_{a}^{2}c_{a}m_{a}^{(e)}=-\frac{4Iq^{2}\kappa^{2}}{3\eta}\int\frac{d\mathbb{k}}{(2\pi)^{3}}\frac{|\varrho_{\mathbb{k}}|^{2}}{\mathbb{k}^{2}(\kappa^{2}|\varrho_{\mathbb{k}}|^{2}+\mathbb{k}^{2})},$ (57)
where $I=\sum_{a}z_{a}^{2}c_{a}/2$ is the ionic strength.
V Kohlrausch’s law and beyond
Based on the above results, we can write the total conductivity as
$\sigma=\sigma_{0}+\Delta\sigma_{r}+\Delta\sigma_{e},$ (58)
where $\sigma_{0}={q^{2}}\left(z_{1}^{2}/R_{1}+z_{2}^{2}/R_{2}\right)/(6\pi\eta)$ is the Nernst-Einstien conductivity of very dilute solution, and the second and third terms are determined by eqs. (47) and (57), respectively.
In the case of a very dilute solution ($c_{a}\to 0$), where we can approximate the bulk chemical potentials by the ideal-gas expressions
$\mu_{a}=k_{B}T\ln(c_{a}\lambda_{a}^{3}/\xi_{a}),$ (59)
where $k_{B}$ is the Boltzmann constant, $T$ is the temperature, $\lambda_{a}$ is the thermal wavelength of the ion $a$, $\xi_{a}$ is the internal partition function of ion $a$ (which drops out from final expressions), we obtain for very dilute one-component 1:1 electrolyte solution $J_{11}=J_{22}=k_{B}T/c$, $J_{12}=J_{21}=0$, where $c=c_{1}=c_{2}$. Thus in the case of point-like charge when $\varrho_{\mathbb{k}}=1$ after calculations of integrals, as expected, the developed theory results in the Debye-Hückel-Onsager expression [18, 19, 20] for total conductivity of 1:1 electrolyte solution (so-called Kohlrausch’s law [14, 20])
$\sigma=\sigma_{0}(1-\xi l_{B}^{3/2}c^{1/2}),$ (60)
$\xi=\frac{2\sqrt{\pi}}{3}(\sqrt{2}-1)+\frac{4\sqrt{2\pi}R}{l_{B}},$ (61)
where $l_{B}=q^{2}/(4\pi\varepsilon\varepsilon_{0}k_{B}T)$ is the Bjerrum length.
It is clear that Kohlrausch’s law, which was originally obtained for very dilute solutions [18], does not take into account the excluded volume effects of ions and their short-range specific interactions. To take into account the short-range specific interactions of ions, which should be significant in rather concentrated solutions, we can use the following expression for the free energy density:
$f_{id}=k_{B}T\sum\limits_{a}c_{\alpha}\left(\ln(c_{a}\lambda_{a}^{3}/\xi_{a})-1\right),$ (63)
the second term describes the contribution of steric interactions of the ions within the model of hard sphere mixture in the Percus-Yevick approximation [48]:
Note that in this paper, we will only consider the case of aqueous electrolyte solutions, for which we assume that the parameters of short-range specific ion interactions are zero ($A_{ab}=0$). Therefore, we only take into account the effect of excluded volume (steric) interactions. However, a short-range specific interaction is important for room-temperature ionic liquids [49, 50].
VI Model charge form factor
To calculate the relaxation and electrophoretic contributions to conductivity, we need to specify the charge form factor. We will use the following exponential (Slater-type [51]) distribution function:
$\varrho(\mathbb{r})=\frac{1}{8\pi a^{3}}\exp\left[-{r}/{a}\right]$ (65)
with the following Fourier-image
[…]
where $a$ is the characteristic length scale of charge smearing, which should be order of the size of a solvated ion. As we demonstrate below, this phenomenological distribution function provides a very good approximation of the electrical conductivity of different electrolyte solutions at relatively high concentrations.
where $s=\kappa a$, $u=\kappa_{1}/\kappa$ and
[…]
The latter function has the following asymptotic behavior:
The regime where $s\ll 1$ corresponds to the region of validity of the standard DHO theory, while the new regime where $s\gg 1$ is realized for electrolyte solutions with trivalent cations ($z_{1}=3$), room temperature ionic liquids and molten salts (see discussion below).
$\Delta\sigma_{e}=-\frac{2Iq^{2}\kappa}{3\pi^{2}\eta}\theta(s),$ (70)
where
$\theta(s)=\int\limits_{0}^{\infty}\frac{dx}{1+x^{2}(1+x^{2}s^{2})^{4}}.$ (71)
VII Numerical results
Let us consider aqueous one-component electrolyte solutions with different ratios $z_{1}$:$z_{2}$ and compare their theoretically computed electrical conductivities with available experimental data using a model where ions are treated as hard spheres with smeared charges. Our model accounts for the hydration of ions through their hydration diameters, $d_{a}$, and dependence of the dielectric constant on ion concentrations, $\varepsilon=\varepsilon_{w}-\alpha_{1}c_{1}-\alpha_{2}c_{2}$, where $\varepsilon_{w}$ is the dielectric constant of pure water and $\alpha_{1,2}$ are the dielectric decrements [40, 41, 52] of ions. Additionally, we take into account the dependence of the electrolyte solution’s viscosity on the salt concentration, approximating a variety of available experimental data [53, 54, 55, 56, 57, 58] with a Barthel-like exponential relationship [59, 23].
Tables 1 and 2 present the values of the parameters used in the calculation of electrical conductivity and the fitting parameter – the length scale of ionic charge smearing, $a$ – respectively. Fig. 1 presents approximations of the experimental dependencies [53, 56, 58, 60, 61, 56, 62] of the molar electrical conductivity of various aqueous electrolyte solutions on concentration at $T=298.15$ K using the proposed model. As shown, the theory provides a decent approximation of the experimental data over a relatively wide concentration range, up to nearly 4 M for 1:1 electrolytes, with just one adjustable parameter. Despite the generally acceptable reproduction of the experimental data (see Fig. 1), it is important to note that there are slight deviations from the reference data at high concentrations for most of studied electrolyte solutions, which will be discussed further below.
ion $d/2$ [63], Å $R_{h}$ [63], Å $\alpha$ [64, 65], $M^{-1}$
Na+ 3.58 1.84 8
K+ 3.31 1.25 8
Li+ 3.82 2.38 11
Mg2+ 4.28 3.47 24
La3+ 4.52 3.96 35
Cl- 3.32 1.21 3
Br- 3.30 1.18 0
I- 3.31 1.19 7
$\text{NO}^{3-}$ 3.35 1.29 0
$\text{SO}_{4}^{2-}$ 3.79 2.30 7
electrolyte $a$, Å electrolyte $a$, Å
NaCl 7.41 NaI 8.93
KCl 5.39 $\text{KNO}_{3}$ 2.55
LiCl 7.46 $\text{NaNO}_{3}$ 4.47
MgCl2 10.52 $\text{K}_{2}\text{SO}_{4}$ 3.99
LaCl3 9.38 $\text{Na}_{2}\text{SO}_{4}$ 4.88
KBr 4.46 $\text{Li}_{2}\text{SO}_{4}$ 4.55
NaBr 6.86 $\text{Mg}\text{SO}_{4}$ 5.06
KI 4.79
Figure 1: Molar electrical conductivity as a function of concentration for a set of aqueous electrolyte solutions at $T=298.15$ K. The lines represent the best fit of the experimental data [53, 56, 58, 60, 61] using equations (58), (67) and (70). The dashed line in (1(a)) represents Kohlrausch’s law for LiCl aqueous solution.
If we consider the form factor as an essential characteristic of the ion, it should enable us to calculate electrical conductivity at various temperatures, provided we have obtained its value from preliminary experimental data fit. In Fig. 22(a), we present the results of such an attempt for the aqueous solution of KCl, i.e. we approximated the data at $T=298.15$ K and calculated conductivity with a fixed charge smearing parameter $a$ at different temperatures. While this method can estimate conductivity for temperatures close to the one used in the fit, it misses the values of electrical conductivity at infinite dilution. The only parameters in the expression for $\sigma_{0}$, which can be considered as adjustable ones, are the Stokes radii of the ions, or the reduced radius $R$, introduced above, when dealing with a 1:1 electrolyte like KCl. Thus, an effective way to improve temperature-dependent conductivity calculations is to define there radii as functions of temperature. Although it is not straightforward to find experimental data for the conductivity at infinite dilution for each ion, in the case of KCl we can model both ions as identical due to their similar Stokes radii (see Table 1). By fitting the infinite dilution conductivity of KCl aqueous solution at different temperatures [53], we obviously achieve improved agreement between experimental data and model calculations (see Fig. 22(b)).
Extending this approach to other aqueous electrolyte solutions appears straightforward: by considering the temperature dependence of the electrical conductivity at infinite dilution and assuming a constant Stokes radii ratio ($R_{Na}/R_{Cl}\approx 1.5$), we achieve excellent agreement between theoretical and experimental isotherms for the case of NaCl solution (see Fig. 33(a) and 3(b)).
Figure 2: Molar electrical conductivity isotherms of KCl aqueous solution. The lines represent the theoretical predictions with constant (2(a)) and temperature-dependent (2(b)) Stokes radii. The symbols represent the experimental data [53, 62, 58].
Figure 3: Molar electrical conductivity isotherms of NaCl aqueous solution. The lines represent the theoretical predictions with constant (3(a)) and temperature-dependent (3(b)) Stokes radii. The symbols represent the experimental data [53, 62, 58].
VIII Discussion
As previously noted, the model results show slight deviations from experimental data at high concentrations for most of studied systems. The most significant discrepancies occur for multivalent ions, including sulfates of lithium, sodium, and potassium, not presented in Fig. 1. For these systems, the theory provides a reasonably good approximation only up to a concentration of approximately 0.3 M. Minor contributions to observed deviations can be attributed to potential inaccuracies arising from experimental errors and the process of viscosity and electrical conductivity data fitting, as well as clear uncertainty related to the choice of dielectric decrement values  [64, 65]. However, the primary cause of these deviations is likely due to the two main assumptions made in the utilized model. Firstly, the model neglects the formation of ionic pairs, which could significantly impact the results at rather high concentrations. Indeed, for monovalent electrolytes, ion pairs in a solution can form electrically neutral particles, which renormalize the total dielectric constant, whereas in the case of multivalent ones, ion pairs lead to the formation of ions with reduced charge. Thus, the contribution of excluded volume interactions cannot be accurately described within the hard spheres model used in this study. It may be more appropriate to represent ion pairs as dumbbells [67, 68]. Using the expression for the total free energy of a mixture of convex bodies [69] may provide a more accurate description of the steric interactions between these particles. Secondly, the model assumes identical form factors for both ions of the electrolyte, which may be an oversimplification, particularly for multivalent ions. Additionally, the model has limitation related to the concentration limit up to which conductivity can be accurately calculated. This limitation arises from the equation of state for hard spheres. Specifically, the ion concentrations must satisfy the condition $\pi\sum_{a}c_{a}d_{a}^{3}/6<1$ (see equation (64)). For example, the maximum concentration at which this theory is applicable is approximately 4.2 M for 1:1 electrolyte solutions. For 2:1, 1:2, and 3:1 electrolytes, the upper concentration limits are even lower. However, this limitation is not critical, as the developed model covers a sufficiently wide range of concentrations for all practically significant aqueous electrolyte solutions (see Fig. 1).
Let us now compare the magnitudes of the electrophoretic and relaxation contributions to total conductivity and explore how they vary with changes in concentration and temperature. As shown in Fig. 4, the electrophoretic contribution is significantly larger than the relaxation one. This aligns with recent studies, which successfully described the conductivity of aqueous electrolyte solutions primarily by considering the electrophoretic effect [28, 27]. The relaxation correction initially increases at low concentrations, peaks, and then declines towards zero. Small values of relaxation correction in concentrated solutions can be attributed to the strong binding of ions with their surroundings, which minimizes the distortion of the ionic atmosphere in the presence of an external electric field. In contrast, the electrophoretic correction steadily increases with concentration because a higher number of ions in the solution generates a stronger volume force on the solvent in an external field. Fig. 5 illustrates that both contributions to electrical conductivity increase with temperature. This is due to the reduction in electrostatic ’binding’ between ions and their ionic atmospheres, which makes it easier for the ionic atmosphere to be distorted and carried away by an external field.
Figure 4: The absolute values of the electrophoretic (red line) and relaxation (blue line) contributions to the electrical conductivity, as functions of concentration. The values are reduced to the Nernst-Einstein conductivity, $\sigma_{0}$, for a KCl aqueous solution at $T=298.15$ K.
Figure 5: The concentration dependencies of absolute values of the electrophoretic (5(a)) and relaxation (5(b)) contributions to molar electrical conductivity for the case of KCl aqueous solution, plotted at different temperatures.
To obtain a precise estimate of the experimental data on electrical conductivity, it is important to consider the impact of electrolyte additives on the shear viscosity and dielectric constant of the solution. The literature provides this information for most electrolytes relevant to our research. Although neglecting these effects may affect accuracy, using the values of the pure solvent for these properties can still yield reasonably accurate estimates, as demonstrated in recent studies [27, 28, 70]. It is important to note that the fits of the experimental data on electrical conductivity as a function of concentration were performed for systems with ions whose dielectric decrements are known. However, this is not a fundamental limitation, as empirical formulas can still be used in cases where data on ion decrements are unavailable [23] to describe the dependence of the dielectric constant of the aqueous electrolyte on concentration. In a similar manner, empirical formulas can also be applied to estimate the shear viscosity of aqueous electrolyte solutions.
It is interesting to note that asymptotic behavior of $\kappa a\gg 1$ (eqs. (69) and (72)) can be achieved for 3:1 aqueous electrolyte solutions at sufficiently high concentrations. As shown in Fig. 6, this regime for LaCl3 aqueous solutions occurs at concentrations of approximately 1-1.5 M. This regime can also be achieved for room-temperature ionic liquids and molten salts. For room-temperature ionic liquids, it is more relevant to use the previously mentioned equation of state for mixtures of convex bodies, such as dumbbells [67], to account for the specific shape of molecular ions (e.g., the length of the alkyl chain of cations). Additionally, the formation of ionic clusters in room-temperature ionic liquids should be considered, as it can reduce the total electrical conductivity [71, 70, 72]. Furthermore, on the basis of general considerations, approximating the unique charge form factor in the case of room-temperature ionic liquids would be quite rough, and defining these form factors would be a separate task. Finally, it would be interesting to explore the potential of this theory for modeling the electrical conductivity of solutions containing room-temperature ionic liquids [73, 74]. These topics could be the focus of future research.
IX Conclusion
The nonlocal electrostatics-based mean field theory of electrolyte solutions, which is introduced in this paper as an extension of the classical Debye-Hückel-Onsager theory, takes into account the specific characteristics of ions, such as their steric interactions, hydration, and spatial charge distributions. This model provides analytical expressions for the mobilities of ions and the electrical conductivity, while also considering relaxation and hydrodynamic effects. At low concentrations, the theory agrees well with Kohlrausch’s law, as expected. At higher concentrations, however, it shows good agreement with experimental data that deviate significantly from Kohlrausch’s limiting law. The deviation from Kohlrausch’s law can be attributed to various physical and chemical factors, including steric effects, ion hydration, and nonlocal charge distribution in solvated ions. By introducing a single adjustable parameter to describe the scale of the nonlocal charge distribution of solvated ions, and using experimental data on shear viscosity, Stokes radii, hydration radii, and dielectric decrements of ions, the model was able to successfully approximate experimental data on electrical conductivity of aqueous electrolyte solutions over a wide concentration range. These results provide predictive insights that are consistent with existing experimental findings. The theory can be readily expanded to estimate the electrical conductivity of binary electrolyte solutions for which there is a vast amount of accessible experimental data.
Acknowledgements. YAB would like to express its gratitude to Olga I. Vinogradova for bringing to his attention some issues raised in the paper. The calculations were performed using computational resources from HPC facilities at NRU HSE. The article was prepared within the framework of the project ’Mirror Laboratories’ HSE University. NNK and YAB would like to thank L.P. Safonova, A.M. Kolker, L.E. Shmukler, M.A. Krestyaninov, and D.M. Makarov for their assistance in finding the experimental data. NNK also appreciates S.E. Kruchinin’s assistance with the numerical computations.
@@ A. Szcz
  • Landau and Lifshitz [2013] L. D. Landau and E. M. Lifshitz, Fluid mechanics: Landau And Lifshitz: course of theoretical physics, Volume 6, Vol. 6 (Elsevier, 2013).

### v2

@@ IX Conclusion
$m_{a}^{(e)}=-\frac{2\kappa^{2}}{3\eta}\int\frac{\text{d}\mathbb{k}}{(2\pi)^{3}}\frac{|\varrho_{\mathbb{k}}|^{2}}{\mathbb{k}^{2}(\mathbb{k}^{2}+\kappa^{2}|\varrho_{\mathbb{k}}|^{2})}.$ (74)
Thus, taking into account that $\Delta\sigma_{e}=\sum_{a}q_{a}^{2}c_{a}m_{a}^{(e)}$, we immediately obtain eq. (41).
Appendix C
electrolyte RAAD, % $c_{max},\leavevmode\nobreak\ mol\cdot l^{-1}$ $N$
KCl 1.34 4 160
NaCl 1.71 4 115
LiCl 1.62 4 42
CaCl2 0.94 1.9 31
MgCl2 2.46 1.9 53
LaCl3 0.93 1.4 45
KBr 0.62 3.8 31
NaBr 1.34 4 24
KI 0.88 4 74
NaI 2.70 4 33
$\text{KNO}_{3}$ 0.98 3 33
$\text{NaNO}_{3}$ 1.20 4 32
$\text{LiNO}_{3}$ 2.97 3.7 10
$\text{K}_{2}\text{SO}_{4}$ 11.74 0.65 19
$\text{Na}_{2}\text{SO}_{4}$ 15.51 1.4 25
$\text{Li}_{2}\text{SO}_{4}$ 19.48 2.5 25
$\text{Mg}\text{SO}_{4}$ 26.44 2 28
Table 3: Relative Absolute Average Deviation (RAAD) of studied aqueous electrolyte solutions at 298.15 K; $c_{max}$ is the highest concentration within the studied range, $N$ is a number of experimental points, RAAD$=\frac{1}{N}\sum\left(\frac{|\sigma_{exp}-\sigma_{calc}|}{\sigma_{exp}}\right)\times 100$.
Figure 9: Linear correlation between the ion charge smearing length, $a$, and the mean square hydration radius, $\sqrt{(r_{+}^{2}+r_{-}^{2})/2}$, of the ions for the chlorides at 298.15 K. Coefficient of determination $\approx 0.91$.
@@ A. Szcz
  • Fraenkel [2018] D. Fraenkel, ‘‘An improved theory of the electric conductance of ionic solutions based on the concept of the ion-atmosphere’s smaller-ion shell,’’ Physical Chemistry Chemical Physics 20, 29896–29909 (2018).
  • Fraenkel [2010] D. Fraenkel, ‘‘Simplified electrostatic model for the thermodynamic excess potentials of binary strong electrolyte solutions with size-dissimilar ions,’’ Molecular Physics 108, 1435–1466 (2010).
  • Henry [1931] D. Henry, ‘‘The cataphoresis of suspended particles. part i. - the equation of cataphoresis,’’ Proceedings of the Royal Society of London. Series A, Containing Papers of a Mathematical and Physical Character 133, 106–129 (1931).
@@
  • [45] D. Mazur, P. Brandyshev, S. Doronin, and Y. A. Budkov, ‘‘Understanding the electric double layer at the electrode-electrolyte interface: Part i-no ion specific adsorption,’’ ChemPhysChem , e202400650.
@@
  • Landau and Lifshitz [2013] L. D. Landau and E. M. Lifshitz, Fluid mechanics: Landau And Lifshitz: course of theoretical physics, Vol. 6 (Elsevier, 2013).
@@
  • Posel et al. [2014] Z. Posel, Z. Limpouchova, K. Sindelka, M. Lisal, and K. Prochazka, ‘‘Dissipative particle dynamics study of the ph-dependent behavior of poly (2-vinylpyridine)-block-poly (ethylene oxide) diblock copolymer in aqueous buffers,’’ Macromolecules 47, 2503–2514 (2014).
  • Mao et al. [2015] R. Mao, M.-T. Lee, A. Vishnyakov, and A. V. Neimark, ‘‘Modeling aggregation of ionic surfactants using a smeared charge approximation in dissipative particle dynamics simulations,’’ The Journal of Physical Chemistry B 119, 11673–11683 (2015).
  • Hendrikse, Amador, and Wilson [2024] R. L. Hendrikse, C. Amador, and M. R. Wilson, ‘‘Dpd simulations of anionic surfactant micelles: a critical role for polarisable water models,’’ Soft Matter  (2024).
@@
  • JThG [1943] O. JThG, ‘‘Theorie der elektrophorese,’’ Kolloid-Beihefte 54, 287 (1943).